The perimeter of a rectangle is 48 feet. One dimension is
16 feet. Find the area of the rectangle.

Answer by Cromlix(4381)

Perimeter = 2*length + 2*width
48 = 2*16 + 2*width
48 = 32 + 2width
48 -32 = 2 width
16/2 = width
Width = 8 feet
Area = 8 * 16 = 128 feet^2
Hope this helps.
:-)